The massless spring of a spring gun has a force constant k = 0.04 N/cm. When the gun is aimed vertically, a 30 g projectile is shot to a height of 5.0 m above the end of the expanded spring. How much was the spring compressed initially? (Find d, in meters.)
1
Expert's answer
2019-03-07T11:03:40-0500

In accordance with the law of conservation of energy kd2/2 = mgh , then d = (2mgh/k)1/2 = (2*0.3*5/4)1/2 = 0.87 (m).


d = 0.87 m